Temperature is measured in two scales (there are others too): Fahrenheit and Centigrade. The equation that converts C degrees to F degrees is:
f=(9/5)c+2
Now, take the above equation and solve it for C so you can convert the other way (convert F degrees to C degrees). How would I do this?

Answer by guapa(62)

First of all: The formula you stated is not quite right. F=(9/5)C+32 (not +2)
Now, solving this formula for C you have to first divide both sides by 9/5. The rule says that dividing by a fraction you multiply by its reciprocal which is 5/9. Thus:
F*(5/9)= C+32*(9/5)*(5/9)
F(5/9)= C+32 Now substract 32 from both sides
F-32(5/9) = C+32-32
Your final formula is:
C=(5/9)F-32